Walt made an extra $5000 last year from a part-time job. He invested part of the money at 6% and the rest at 10%. He made a total of $420 in interest. How much was invested at 10%?
Answer by tutorcecilia(2152)

Walt made an extra $5000 last year from a part-time job. He invested part of the money at 6% and the rest at 10%. He made a total of $420 in interest. How much was invested at 10%?
Use: x+y=c
So, x+y=5000
y=50000-x
and
Use: ax+by=total
So, .06x+.10y=420
.06x+.10(5000-x)=420 [substitute y=5000-x)
.06x+500-.10x=420
.06x-.10x=420-500
-.04x=-80
-0.4x/.04=-80/-.04
x=2000
Substituting: 2000+y=5000
y=3000
answer: y=$3000